OpenStudy (anonymous):

Algebra Two Help: What is the solution of log(2x − 3)125 = 3 I'm not sure how to properly type logs but the exponential form is (2x-3)^3=125.

OpenStudy (anonymous):

The goal is to solve for X

OpenStudy (anonymous):

x = 4

OpenStudy (anonymous):

(2x-3)^3=5^3 2x-3=5 2x=5+3 x=?